matthew is 8 years older than twice daniel's age. the sum of twice matthew's age and four times daniel's age is 176. how old are matthew and daniel
Answer by Cromlix(4381)

Hi there,
matthew is 8 years older than twice daniel's age
M = 8 + 2D
the sum of twice matthew's age and four times
daniel's age is 176.
2M + 4D = 176
Replace M with (8 + 2D) from top equation.
2(8 + 2D) + 4D = 176
16 + 4D + 4D = 176
8D = 176 - 16
8D = 160
D = 20
Daniel is 20 years old
Matthew is 48 years old
Hope this helps :-)